Exercise 25: Let f : [a, b] → R be continuous and ϕ : [a, b] → R be integrable and
non-negative, i.e. ϕ ≥ 0. Then there exists c ∈ (a, b) such that
the integral of f(x)Ï•(x)dx from a to b = f(c) the integral of Ï•(x)dx from a to b.
1
Expert's answer
2020-11-19T12:53:43-0500

Let "m=\min\limits_{x\in[a,b]}f(x)" and "M=\max\limits_{x\in[a,b]}f(x)".

Then "\varphi(x)(f(x)-m)\ge 0" on "[a,b]" , so "\int\limits_a^b \varphi(x)(f(x)-m)dx\ge 0", that is "m\int\limits_a^b\varphi(x)dx\le\int\limits_a^b \varphi(x)f(x)dx".

Similarly we can obtain "\int\limits_a^b \varphi(x)f(x)dx\le M\int\limits_a^b \varphi(x)dx"

So we have 2 cases:

1)If "\int\limits_a^b \varphi(x)dx=0", then "\int\limits_a^b \varphi(x)f(x)dx=0=f(c)\int\limits_a^b \varphi(x)dx" for every "c\in(a,b)"

2)If "\int\limits_a^b \varphi(x)dx\neq 0", then "m\le\frac{\int\limits_a^b \varphi(x)f(x)dx}{\int\limits_a^b \varphi(x)dx}\le M". Then by the intermediate value theorem there is "c\in[a,b]" such that "\frac{\int\limits_a^b \varphi(x)f(x)dx}{\int\limits_a^b \varphi(x)dx}=f(c)", that is "\int\limits_a^b \varphi(x)f(x)dx=f(c)\int\limits_a^b \varphi(x)dx".

Prove that if "c\in\{a,b\}", then there is "c'\in(a,b)" such that "\int\limits_a^b \varphi(x)f(x)dx=f(c')\int\limits_a^b \varphi(x)dx". We need to consider 2 subcases:

2.1)"c=a". Then we have "\int\limits_a^b \varphi(x)(f(x)-f(a))dx=0".

If there is "c'\in(a,b)" such that "f(a)=f(c')", then "\int\limits_a^b \varphi(x)f(x)dx=f(c')\int\limits_a^b \varphi(x)dx"

Suppose that there is not "x\in(a,b)" such that "f(x)=f(a)", then by the intermediate value theorem "f(x)-f(a)" does not change the sign on "(a,b)". Without loss of generality we can assume that "f(x)>f(a)" for every "x\in(a,b)".

So we have that "\varphi(x)(f(x)-f(a))\ge 0" on "(a,b)". Then since "\int\limits_a^b \varphi(x)(f(x)-f(a))dx=0", we have "\varphi(x)(f(x)-f(a))=0" almost everywhere on "(a,b)". Since "f(x)-f(a)>0" on "(a,b)", we have that "\varphi(x)=0" almost everywhere on "(a,b)", that is "\int\limits_a^b \varphi(x)dx=0".

By the point 1) we obtain that "\int\limits_a^b \varphi(x)f(x)dx=f(c')\int\limits_a^b \varphi(x)dx" for every "c'\in(a,b)"

So we obtain that there is "c'\in(a,b)" such that "\int\limits_a^b \varphi(x)f(x)dx=f(c')\int\limits_a^b \varphi(x)dx"

2.2)"c=b". This case is similar to 2.1.

So for the case 2 we obtain that "\int\limits_a^b \varphi(x)f(x)dx=f(c')\int\limits_a^b \varphi(x)dx" for some "c'\in(a,b)".

Therefore in every case we obtain that "\int\limits_a^b \varphi(x)f(x)dx=f(c)\int\limits_a^b \varphi(x)dx" for some "c\in(a,b)"